OpenStudy (anonymous):

If the Federal Reserve sets the reserve rate to 5%, what is the resulting money multiplier?

OpenStudy (anonymous):

There's a reciprocal relationship between the reserve ratio and the money multiplier. If the rate is \(k\), then the multiplier is \(\dfrac{1}{k}\). Given that \(k=5%=0.05\), you have \(\dfrac{1}{0.05}=\dfrac{1}{\frac{5}{100}}=\dfrac{100}{5}=20\).

OpenStudy (anonymous):

*\(k=0.05\).